# Unveiling the Basin-Like Loss Landscape in Large Language Models

### v2

@@ F.1 Randomized Smoothing Regime
As models grow larger, we draw the following conclusions:
  • As models grow larger, their basins naturally expand, increasing the theoretically guaranteed radius (which is $O(\sigma)$) for subsequent fine-tuning, making it harder to compromise prior capabilities.
  • As models and their basins grow larger, the regions within the basin exhibit greater expressive power, facilitating the acquisition of new capabilities.
We define the region where the model is guaranteed to preserve prior capabilities while learning new capabilities as the Randomized Smoothing Regime (RS Regime). The region where the model is guaranteed to preserve prior capabilities has a size of $O(\sigma)$ (see Sec.˜4.1). Taking Rademacher complexity as a metric for expressive power, for the expressive power to exceed $k$, we require $O(d\sigma^{2})\geq k\iff\sigma\geq O\left(\sqrt{\frac{k}{d}}\right)$. Thus, the RS Regime is $0,O(\sigma)]\cap\left[O\left(\sqrt{\frac{k}{d}}\right),+\infty\right)$. As models grow larger, $O(\sigma)$ continues to grow, while $O\left(\sqrt{\frac{k}{d}}\right)$ continues to decrease. Consequently, the RS Regime, representing the theoretically guaranteed region for preserving prior capabilities and acquiring new capabilities, continuously expands. When models have infinitely many parameters, all fine-tuning may lie within the RS Regime, analogous to the Neural Tangent Kernel (NTK) regime (Jacot et al., 2018), where infinitely wide models behave like kernel methods. Future work could explore how quickly models enter the RS Regime as their size increases and identify when this transition occurs.

### v3

@@ D.10 Landscape of Vision Language Models (VLMs)
E.4 About Norm Constraints in Landscape Visualization
E.5 Theoretical Intuition: Stability via Lazy Training Dynamics
To provide theoretical intuition for our observation that scaling up models facilitates the coexistence of preserving old capabilities and acquiring new ones, we refer to the well-established Lazy Training phenomenon in the Neural Tangent Kernel (NTK) regime (Jacot et al., 2018). Existing analyses in this field demonstrate that when using the NTK parameterization, in the infinite-width limit, the parameter displacement required to learn a new task vanishes relative to the scale of initialization, ensuring the model remains within the local linear region.
Proposition E.2 (Basin Stability in the Infinite-Width Limit (Jacot et al., 2018; Du et al., 2019)).
Consider a neural network $f(\bm{x};\bm{\theta})$ with width $m$ using NTK parameterization. In the limit $m\to\infty$, the parameter update $\Delta\bm{\theta}$ required to minimize the fine-tuning loss $\mathcal{L}_{ft}$ goes to zero.
Derivation from Literature.
The above proposition adapts standard results from NTK literature (e.g., Du et al. (2019); Allen-Zhu et al. (2019)) to the context of fine-tuning stability. We outline the key steps to clarify the dependency between model width and optimization trajectory.
1. Linearization and Minimum Norm Solution. Consider the linearized model $g_{\bm{\theta}}(\bm{x})$ around initialization $\bm{\theta}_{0}$. The training objective for this linear model is a least-squares problem:
$\min_{\Delta\bm{\theta}}\frac{1}{2}\|\Phi\Delta\bm{\theta}-\vec{y}\|_{2}^{2},$ (8)
where $\Phi\in\mathbb{R}^{n\times p}$ is the Jacobian matrix and $\vec{y}$ is the target residual vector. Assuming over-parameterization ($p\geq n$) and a minimum singular value $\sigma_{\min}(\Phi)=\sigma>0$, the minimum-norm solution is given by the Moore-Penrose pseudoinverse $\Delta\hat{\bm{\theta}}=\Phi^{\top}(\Phi\Phi^{\top})^{-1}\vec{y}$. The norm of this update is bounded by:
$\|\Delta\hat{\bm{\theta}}\|_{2}\leq\|\Phi^{+}\|_{\text{op}}\|\vec{y}\|_{2}=\frac{1}{\sigma}\|\vec{y}\|_{2}=O\left(\frac{\sqrt{n}}{\sigma}\right).$ (9)
This derivation clarifies why, in a sufficiently well-conditioned landscape (large $\sigma$), the optimal solution lies within a small ball around the initialization.
2. Approximation via Taylor Expansion. The validity of substituting the non-linear $f$ with the linear $g$ hinges on the smoothness of the loss landscape. Suppose the gradient $\nabla_{\bm{\theta}}f$ is $\beta$-Lipschitz. By the standard descent lemma (a consequence of the second-order Taylor expansion), the difference between the function and its first-order approximation is bounded by the quadratic term of the parameter change:
$|f(\bm{x};\bm{\theta})-g_{\bm{\theta}}(\bm{x})|=\left|f(\bm{\theta})-\left(f(\bm{\theta}_{0})+\nabla f(\bm{\theta}_{0})^{\top}\Delta\bm{\theta}\right)\right|\leq\frac{\beta}{2}\|\Delta\bm{\theta}\|_{2}^{2}.$ (10)
Substituting the bound from Step 1, we obtain the approximation error bound:
$|f(\bm{x};\bm{\theta})-g_{\bm{\theta}}(\bm{x})|\leq\frac{\beta}{2}\cdot O\left(\frac{n}{\sigma^{2}}\right)=O\left(\frac{\beta n}{\sigma^{2}}\right).$ (11)
3. Scaling Behavior. Crucially, random matrix theory establishes that as width $m\to\infty$, the ratio $\frac{\beta}{\sigma^{2}}\propto\frac{1}{\sqrt{m}}\to 0$. Consequently, the approximation error vanishes, and the optimization trajectory of the deep network stays strictly close to that of the linear model.
Implication for Our Findings. These established theoretical results offer an explanation for our empirical observation in Sec.˜3.2: as models scale up, they may naturally enter a regime where fine-tuning induces minimal parameter movement relative to the stability basin, effectively mitigating catastrophic forgetting for benign tasks.
@@ F.1 Randomized Smoothing Regime
As models grow larger, we draw the following conclusions regarding the interplay between robustness and learnability: (1) As models grow larger, their basins naturally expand, increasing the theoretically guaranteed radius (which is $O(\sigma)$) for subsequent fine-tuning, making it harder to compromise prior capabilities. (2) As models and their basins grow larger, the regions within the basin exhibit greater expressive power, facilitating the acquisition of new capabilities.
We define the region where the model is guaranteed to preserve prior capabilities while possessing sufficient capacity to learn new capabilities as the Randomized Smoothing Regime (RS Regime). The region where the model is guaranteed to preserve prior capabilities has a radial size of $O(\sigma)$ (see Sec.˜4.1). Taking Rademacher complexity (Neyshabur et al., 2015) as a metric for expressive power, for the expressive power to exceed a complexity $k$, we require the available radius to satisfy $O(d\sigma^{2})\geq k\iff\sigma\geq O\left(\sqrt{\frac{k}{d}}\right)$. Thus, for a fine-tuning task to be feasible and safe, the modification distance must fall within the intersection $0,O(\sigma)]\cap\left[O\left(\sqrt{\frac{k}{d}}\right),+\infty\right)$.
As models grow larger ($d\to\infty$), the safety bound $O(\sigma)$ continues to grow, while the required learning radius $O\left(\sqrt{\frac{k}{d}}\right)$ continues to decrease. Consequently, the RS Regime—representing the feasible region for simultaneously preserving prior capabilities and acquiring new ones—continuously expands. In the limit, when models have infinitely many parameters, all fine-tuning trajectories may lie within the RS Regime, analogous to the Neural Tangent Kernel (NTK) regime (Jacot et al., 2018), where infinitely wide models behave like kernel methods (See our analysis in Sec.˜E.5). Future work could explore how quickly models enter the RS Regime as their size increases and identify when this phase transition occurs.
Limitation of the RS Regime. While the RS Regime theoretically guarantees the dual objectives of safety and learnability in the asymptotic limit, current finite-sized models do not yet fully reside within this ideal regime. In this pre-asymptotic stage, capability degradation exhibits significant heterogeneity depending on the fine-tuning data distribution—as evidenced in Sec.˜3.4, where i.i.d. (benign), distribution-shifted, and adversarial fine-tuning trajectories diverge significantly in their degradation rates. Consequently, in scenarios constrained by compute or model size—where one cannot simply enlarge the basin (via suppressing Gaussian noise) sufficiently to encompass the entire fine-tuning trajectory—ensuring model safety and robustness requires complementary alignment techniques beyond pure landscape optimization.
@@ F.2 Larger Scale Pre-training Studies
Appendix G LLM Usage
In the preparation of this manuscript, we utilized large language models, solely for sentence-level language polishing to enhance clarity and readability. The LLMs were used to refine the phrasing of existing text, with all outputs manually reviewed and edited by the authors to ensure accuracy and alignment with the intended scientific content. No LLMs were used in the generation of ideas, experimental design, data analysis, or other scientific contributions in this work.